Solve the system.
2X+Y-3Z=0
4x+2y-6Z=0
X-Y+Z=0

Answer by Alan3354(69443)

2X+Y-3Z=0
4x+2y-6Z=0
X-Y+Z=0
------------
x = y = z = 0
==========
There's no unique solution.
The 2nd equation is 2 times the 1st --> dependent.